# Dynamical mass of a solar-like oscillator at the main-sequence turnoff from Gaia astrometry & ground-based spectroscopy

### v1

@@
Dynamical mass of a solar-like oscillator at the main-sequence turnoff from Gaia astrometry & ground-based spectroscopy
Asteroseismology is widely used for precise determining of masses of solar-like oscillating stars by performing individual-frequency modeling or applying homological scaling relations. However, these methods lack dynamical validation on the main sequence due to the absence of eclipsing double-lined binary system (SB2) as benchmark objects. By providing the orbital inclination, astrometric binary systems from ESA Gaia DR3 offer an abundant alternative for eclipsing systems. We present KIC 9693187 as the first SB2, hosting a solar-like oscillating post-main-sequence star with dynamical masses. By combining Gaia astrometry with spectroscopic obtained with the Las Cumbres Observatory network (LCO), we find = 0.99 0.05 and = 0.89 0.04 for the primary and secondary, respectively. Asteroseismic parameters were extracted from photometry of the NASA Kepler satellite. The mass from individual frequency modeling is = 0.92 0.01. Taking into account the systematic uncertainty of 0.04 for best fit models from individual frequency fitting, we find an agreement within 1.2. From scaling relations we obtain a mass range of 0.93 to 0.98 by using the observed large frequency separations () in the scaling relations for the primary. By using standard corrections for departures from the asymptotic regime of , we obtained a mass range of 0.83 to 1.03. The upper ends of both ranges agree well with the dynamical mass of the primary. This approach provides the first empirical validation for main-sequence solar-like oscillators and opens a new window for validating asteroseismology. Through a dedicatded program targeting astrometric SB2 binary systems, ESA’s PLATO space mission will provide will enlarge the benchmark sample substantially.
@@ 1 Introduction
A star’s mass is the fundamental factor governing its structural properties and evolutionary timescales (Serenelli et al., 2021, and references therein). However, it is challenging to determine this fundamental parameter with precision and accuracy. Asteroseismology is a powerful technique for determining stellar mass by analyzing the pattern of oscillation modes in the power spectral density (PSD). The exquisite photometric quality of space telescopes such as the CoRoT (Convection, Rotation & planetary Transits, Baglin et al., 2009), NASA Kepler, K2 (Borucki et al., 2010; Howell et al., 2014, resp.) or TESS (Transiting Exoplanet Survey Satellite, Ricker et al., 2014) missions have provided data for several 105 of oscillating stars (García and Ballot, 2019).
Stellar masses may be determined by exploiting the power excess of convectively excited solar-like oscillating stars, with two approaches being most commonly used. If the PSD contains a sufficient number of high-quality oscillation modes with identified spherical degrees (), a best-fitting stellar model can be searched for by modeling the individual frequencies (Lebreton and Goupil, 2014; Mathur et al., 2012; Metcalfe et al., 2014; Serenelli et al., 2017; Creevey et al., 2017; Li et al., 2024; Buldgen et al., 2025). The stellar radius (), mass (), and age inferred from the best-fitting model are typically reported with uncertainties of 1%, 4%, and 11%, respectively (e.g., Cunha et al., 2021). However, detailed individual frequency modeling, requires high fidelity mode characterization and significant computational resources, that increase significantly the more evolved a star is. Therefore, the second, global approach for seismic parameter determination, becomes relevant: asteroseismic scaling relations.
@@
provide an estimates of the mass and radius of the solar-like oscillators in solar units (see also Stello et al., 2009; Sharma et al., 2016; Hekker et al., 2020; Belkacem et al., 2013). These equations use the frequency of oscillatory power excess (), large-frequency separation () between modes of identical , consecutive radial orders (), and effective temperature () measured from the star and the Sun () as input for these homological relations. The factor corrects for deviations of the observed value of . These relations are widely applied on large samples in galactic archaeology, stellar evolution and exoplanet host characterization (e.g. Casagrande et al., 2016; Mathur et al., 2016; Pinsonneault et al., 2025; Huber et al., 2022). Typically a precision of 3% in radius, and 6% in mass are reported from large sample studies (e.g., Serenelli et al., 2017). However, both methods still lack robust comparison with model-independent mass estimates, particularly on the MS.
@@
The only direct test for seismic masses on the MS was performed by Gaulme et al. (2016b), by studying the actual oscillation of the Sun, using its reflection on Neptune from K2 photometry. The authors found the solar mass and radius overestimated by 14% and 4%, respectively from scaling relations. They attributed this to shifts in frequency due to enhanced solar activity. Indeed, the best-fit model constrained through individual frequencies led to a mass and radius consistent with simultaneous observations from the ESA Solar and Heliospheric Observatory (SOHO, Domingo et al., 1995). The lack of benchmark systems has left a critical gap in validating seismic masses precisely where they are most often assumed to be robust.
@@
Eclipsing SB2 systems provide model-independent measurements of the stellar masses, constrained from the orbital motion (e.g. Prša, 2018). Comparisons of the dynamical with the seismically inferred mass for red giants have suggested that asteroseismology could overestimate the stellar mass by up to 15% (Gaulme et al., 2016a). Such dichotomy between seismic and dynamical masses was further discussed by several other papers (e.g. Benbakoura et al., 2021; Li et al., 2022). Despite large space-photometry datasets, the benchmark sample of eclipsing SB2 systems with solar-like oscillators remains small (20) and is heavily biased toward H-shell burning red giants (Beck et al., 2024). While 30% of all known main-sequence (MS) solar-like oscillators are in binary systems (Beck, 2025), only few binaries have been seismically studied (Kjeldsen et al., 2005; White et al., 2017; Metcalfe et al., 2012; Beck et al., 2017; García et al., 2025).
@@
The key challenge in the MS mass regime (0.8 1.6, K5-F5) has been the lack of systems with known orbital inclinations, which are necessary to derive dynamical masses. Eclipses naturally provide this constraint from their light curve models. However, the third data release (DR3; Gaia Collaboration et al., 2023) from ESA’s Gaia mission (Gaia Collab. et al., 2016) has opened a new avenue. For a subset of systems, the Gaia non-single star catalog (NSS, Gaia Collab. et al., 2023) provides orbital inclinations for astrometric and single-lined binaries (ASB1). While the inclination is normally the bottleneck for the determination of dynamical masses, for the case of astrometric systems this parameter is known. Thus, the missing piece becomes the detection of the spectroscopic signature of the secondary component from ground-based observations.
@@
In this Letter, we report on the G1V star KIC 9693187111= Gaia DR3 2107491287660520832 = 2MASS J18510009+4625209, the least evolved solar-like oscillator so far, for which dynamical masses can be derived through a combination of Gaia DR3 astrometry and ground-based SB2 spectroscopy from the Las Cumbres Observatory Global Telescope Network (LCO, Brown et al., 2013). Chaplin et al. (2014) identified and characterized as a solar-like oscillator, and the color-magnitude diagram suggests a position on the MS or an early sub-giant (eSG) (Godoy-Rivera et al., 2025). Molenda-Żakowicz et al. (2013) reported this system as an SB2 system, but did not provide orbital parameters. Eventually, the Gaia Collab. et al. (2023) reported this system as ASB1, in the Two-Body Orbit solutions (TBO), within the NSS. In addition to the orbital parameters (=103.750.13 days, =0.390.04), the TBO also reports the inclination angle, of the orbital plane to be 662 degrees.
@@ 2 Spectroscopic analysis and dynamical masses
We monitored KIC 9693187 with NRES (Network of Robotic Echelle Spectrographs, Brown et al., 2013), mounted on the 1 m telescopes of the northern LCO nodes at McDonald Observatory (USA) and Wise Observatory (Israel). NRES is an échelle spectrograph with a resolution of 53,000. Between April 2023 and June 2024, we have collected 22 spectra of KIC 9693187.
The NRES pipeline calculates the cross-correlation function’s response profile (CCF) from the observed spectra and a synthetic spectrum as a template to compute the radial velocitis (RV). The pipeline automatically selected the most appropriate template from the Phoenix model library, based on the Gaia parameters of the target ( = 5700 K, = 4 dex, [Fe/H] = 0 dex). The CCF profiles of the individal spectra, shown in Fig. 4 clearly reveal the SB2 nature of KIC 9693187.
@@ 2.1 Spectral disentangling
To determine the RV semi-amplitudes of the primary and secondary, and respectively, we applied spectral disentangling (SPD) in Fourier space. This technique optimizes the orbital elements of a binary, including and (Simon and Sturm, 1994; Hadrava, 1995). We note that we refer to the more massive component as the primary, indicated with the subscript 1.
For our SPD analysis of KIC 9693187, we use the FDBinary code (Ilijic et al., 2004). We focused on a 60 nm wide region between 476.1 and 536.1 nm, including the Mg-triplet, which is well suited to determine for both components. The disentangled spectra are depicted in Fig. 4. Except for the orbital period, which we fix to the astrometric solution from the Gaia TBO, we treated all orbital parameters as free parameters. The resulting values are reported in Table 2. The uncertainties were derived from a Gaussian fit to the parameter distributions from 5000 bootstrap simulations (see Pavlovski et al., 2023).
The best solution finds =26.298 0.058 km/s and =29.394 0.066 km/s as the RVs semiamplitudes for the primary and secondary component. These values reveal a mass ratio of = =0.8947 0.0028 (11% difference in mass).
@@ 2.2 Stellar fundamental parameters
Fundamental stellar parameters were derived through a spectroscopic analysis using an updated version of the 1D/LTE code BACCHUS (Masseron et al., 2016; Hayes et al., 2022). The full analysis and derived parameters are provided in App. A.1 and Table 2. We find an effective temperature of 5738 84K and a sub-solar metallicity ([Fe/H]=-0.36 0.15dex). The significant enrichment of elements ([/Fe]=+0.25 0.05dex) indicates that the star is member of the older thick-disk population. The projected surface rotation velocity, of the primary and secondary (6.71.5, and 8.02.0 km/s, respectively) are more than three times the current solar value of 1.96 km/s (Beck and Giles, 2005) and about six times faster than the expected at the MS turnoff of the Sun. Even more puzzling, no photometric signature of spot-modulation is seen in the lightcurve or PSD (García et al., 2014; Santos et al., 2021, see Fig. 2 below 1 Hz).
@@ 2.3 Dynamical masses
The dynamical masses for both components are given by,
From the determined orbital parameters and from spectroscopy (Table 2), and and inclination, , provided by Gaia, we obtain =0.9930.046 M⊙ and =0.8890.041 M⊙.
@@ 3.1 Global seismology
To test the asteroseismic scaling relations for the mass (Eq. 2), we use effective temperature and the global seismic parameters of the oscillating component, obtained with the A2Z pipleline and reported by Mathur et al. (2022). Furthermore, we re-extracted the light-curve and calculated the PSD (for details see App. B.1 and Fig. 6), and redetermined the values using the universal pattern module of the apollinaire peakbagging code (Breton et al., 2021), as described in App. B.2. The values from both methods agree within the uncertainty of the respective parameters (see Table 3). However, apollinaire leads to smaller uncertainties and a lower value for .
@@ 3.2 Individual-frequency seismology
Using apollinaire, we extracted 28 frequencies of the spherical degree, = 0, 1, and 2. The extracted frequencies (see Table 5) are depicted in the échelle diagram (Fig. 2). We searched for the best stellar model, constrained by using individual-frequency (IF) seismology. The structural modeling was done with the MESA stellar evolution code package (Modules for Experiments in Stellar Astrophysics, Jermyn et al., 2023; Paxton et al., 2011, and references therein). The frequencies were calculated from each model using the Aarhus adiabatic pulsation codes (Christensen-Dalsgaard, 2008). The modeling setup, and input parameters are described in detail in App. B.4. The model frequencies in the échelle diagram (Fig. 2) present a good fit. The model that fits best the observational constraints is found with =0.920.01 M⊙ and =1.380.01 R⊙ (see Table 3).
4 Discussion & Conclusions
Figure 2 and compares the primary’s dynamical mass with seismically inferred masses from scaling relations and individual frequency modeling (Table 3). From scaling relations, we obtain a mass range of 0.930.03 to 0.980.18 by using the observed from apollinaire and A2Z in the scaling relations for the primary. While the masses from the A2Z values agree within 2%, they suffer from large uncertainties. The apollinaire values lead to a mass 6% below but still within the uncertainties of the dynamical mass.
The large-frequency separation is known to depart from the asymptotic regime increasingly with decreasing radial order. To correct the obtained masses, we test two well-established formalisms from the literature (for details see App. B.2). The corrections proposed by Mosser et al. (2013) appear to underestimate the seismic mass for this late MS target by 23%, based on the apollinaire set of parameters. The best agreement (2%) is found from the correction of Li et al. (2022).
@@ 4 Discussion & Conclusions
The best-fit IF model yields an age of 11.20 0.55 Gyr for the primary component of KIC 9693187 (Table 3). This result confirms the evolved nature of the star, placing it at the main-sequence turnoff, at the point of core hydrogen exhaustion, while the secondary remains on the main sequence (Fig. 8). This evolutionary configuration is consistent with the system’s -enhanced chemical composition. However, given the primary’s advanced age, the rapid surface rotation remains puzzling.
The uncertainty on the derived dynamical masses remains relatively large compared to the sub-percent precision that is in principle achievable with this method (Torres et al., 2010). A major contributor is the relatively large error of on the orbital inclination, which significantly contributes to the uncertainty. Looking ahead, larger sample sizes with improved constraints on orbital inclination and the availability of epoch RV in the forthcoming Gaia DR4 that can be combined with ground based RV monitoring, will enable rapid progress in refining dynamical mass measurements for solar-like oscillators in binaries.
As suggested by Beck et al. (2024), the combination of Gaia astrometry of non-eclipsing systems with SB2 solutions from ground-based spectroscopy, or from forthcoming Gaia DRs, offers a promising and abundant new source of benchmark calibrators. These developments are particularly timely in light of the upcoming ESA mission PLATO (PLAnetary Transits and Oscillations of stars; Rauer et al., 2025), which will provide high-precision asteroseismic data for ten thousands of solar-like oscillators. A significant number of ASB1 systems identified in Gaia DR3 as astrometric binaries and potentially hosting a solar-like oscillator on the MS, SG, or red-giant phase have been selected by the authors as part of the Science Calibration and Validation PLATO Input Catalog (scvPIC). The NRES/LCO monitoring presented in this letter is part of a dedicated follow-up program to build an SB2 sample of such high-value targets. This growing number of well-calibrated benchmark stars enables the broader validation of global asteroseismic methods, and offers a rich ensemble of well calibrated stars to test intricacies of stellar evolution from physics-informed models (e.g. Grossmann et al., 2025; Thomsen et al., 2025; Schimak et al., 2026). Ultimately, these efforts will support PLATO’s core scientific objective of improving the precision of stellar age determinations.

### v2

# Dynamical mass of a solar-like oscillator at the main sequence turnoff from Gaia astrometry & ground-based spectroscopy

@@
Dynamical mass of a solar-like oscillator at the main sequence turnoff from Gaia astrometry & ground-based spectroscopy
Asteroseismology is widely used for the precise mass determination of solar-like oscillating stars, based on individual frequency modeling or homological scaling relations. However, these methods have not been dynamically validated on the main sequence (MS) due to the absence of eclipsing double-lined binary system (SB2) as benchmark objects. By providing the orbital inclination, astrometric binary systems from ESA Gaia DR3 offer an abundant alternative for eclipsing systems. We present KIC 9693187 as the first SB2 hosting a solar-like oscillating post-MS star with dynamical masses. By combining Gaia astrometry with spectroscopic data obtained with the Las Cumbres Observatory network (LCO), we found = 0.99 0.05 and = 0.89 0.04 for the primary and secondary, respectively. The asteroseismic parameters were extracted from photometry of the NASA Kepler satellite. The mass from individual frequency modeling is = 0.92 0.01. Taking into account the systematic uncertainty of 0.04 for best-fit models from individual frequency fitting, we found an agreement within 1.2. From the scaling relations, we obtained a mass range of 0.93 to 0.98 by using the observed large frequency separations () in the scaling relations for the primary. By using standard corrections for departures from the asymptotic regime of , we obtained a mass range of 0.83 to 1.03. The upper ends of both ranges agree well with the dynamical mass of the primary. This approach provides the first empirical validation for MS solar-like oscillators and opens a new window for validating the asteroseismology. Through a dedicated program targeting astrometric SB2 binary systems, ESA’s PLATO space mission will effectively enlarge the benchmark sample to a considerable extent.
@@ 1 Introduction
A star’s mass is the fundamental factor governing its structural properties and evolutionary timescales (Serenelli et al., 2021, and references therein). However, it is challenging to determine this fundamental parameter with precision and accuracy. Asteroseismology is a powerful technique for determining stellar mass by analyzing the pattern of oscillation modes in the power spectral density (PSD). The exquisite photometric quality of space telescopes such as the Convection, Rotation & planetary Transits (CoRoT, Baglin et al., 2009), NASA’s Kepler and K2 (Borucki et al., 2010; Howell et al., 2014, resp.), and Transiting Exoplanet Survey Satellite (TESS, Ricker et al., 2014) missions have provided valuable data on hundreds of thousands of oscillating stars (García and Ballot, 2019).
Stellar masses can be determined by exploiting the power excess of convectively excited solar-like oscillating stars and there are two approaches that are most commonly used for this purpose. If the PSD contains a sufficient number of high-quality oscillation modes with identified spherical degrees (), we can search for the best-fitting stellar model by modeling the individual frequencies (Lebreton and Goupil, 2014; Mathur et al., 2012; Metcalfe et al., 2014; Serenelli et al., 2017; Creevey et al., 2017; Li et al., 2024; Buldgen et al., 2025). The stellar radius (), mass (), and age inferred from the best-fitting model are typically reported with uncertainties of 1%, 4%, and 11%, respectively (e.g., Cunha et al., 2021). However, detailed individual frequency modeling requires high-fidelity mode characterization, requiring significant computational resources, which increase in line with the star’s progressive stage of evolution. Therefore, the second, global approach for seismic parameter determination becomes relevant: asteroseismic scaling relations.
@@
provide estimates of the mass and radius of the solar-like oscillators in solar units (see also Stello et al., 2009; Sharma et al., 2016; Hekker et al., 2020; Belkacem et al., 2013). These equations use the frequency of oscillatory power excess (), large-frequency separation () between modes of identical , consecutive radial orders (), and effective temperature () measured from the star and the Sun () as input for these homological relations. The factor, , corrects for deviations of the observed value of . These relations are widely applied on large samples in galactic archaeology, stellar evolution, and exoplanet host characterization (e.g., Casagrande et al., 2016; Mathur et al., 2016; Pinsonneault et al., 2025; Huber et al., 2022). Typically a precision of 3% in radius and 6% in mass are reported from large sample studies (e.g., Serenelli et al., 2017). However, both methods still lack robust comparison with model-independent mass estimates, particularly on the MS.
@@
The only direct test for seismic masses on the MS was performed by Gaulme et al. (2016b), by studying the actual oscillation of the Sun, using its reflection on Neptune from K2 photometry. The authors found the solar mass and radius to be overestimated by 14% and 4%, respectively, from the scaling relations. They attributed this discrepancy to shifts in frequency due to enhanced solar activity. Indeed, the best-fit model constrained through individual frequencies led to a mass and radius consistent with simultaneous observations from the ESA Solar and Heliospheric Observatory (SOHO, Domingo et al., 1995). The lack of benchmark systems has left a critical gap in validating seismic masses precisely where they are most often assumed to be robust.
@@
Eclipsing SB2 systems provide model-independent measurements of the stellar masses, constrained from the orbital motion (e.g., Prša, 2018). Comparisons of the dynamical mass with the seismically inferred mass for red giants have suggested that asteroseismology could end up overestimating the stellar mass by up to 15% (Gaulme et al., 2016a). Such a dichotomy between the seismic and dynamical masses has been further discussed by several other papers (e.g., Benbakoura et al., 2021; Li et al., 2022). Despite large space-photometry datasets, the benchmark sample of eclipsing SB2 systems with solar-like oscillators remains small (20) and is heavily biased toward H-shell burning red giants (Beck et al., 2024). While 30% of all known (MS) solar-like oscillators are in binary systems (Beck, 2025), only few binaries have been seismically studied (Kjeldsen et al., 2005; White et al., 2017; Metcalfe et al., 2012; Beck et al., 2017)
@@
The key challenge in the MS mass regime (0.8 1.6, K5-F5) has been the lack of systems with known orbital inclinations, which are necessary to derive dynamical masses. Eclipses naturally provide this constraint from their light curve models. However, the third data release (DR3; Gaia Collaboration et al., 2023) from ESA’s Gaia mission (Gaia Collab. et al., 2016) has opened a new avenue. For a subset of systems, the Gaia non-single star catalog (NSS, Gaia Collab. et al., 2023) provides orbital inclinations for astrometric and single-lined binaries (ASB1). While the inclination is normally the bottleneck for the determination of dynamical masses, for the case of astrometric systems, this parameter is known. Thus, the missing piece becomes the detection of the spectroscopic signature of the secondary component from ground-based observations.
@@
In this Letter, we report on the G1V star KIC 9693187111= Gaia DR3 2107491287660520832 = 2MASS J18510009+4625209, the least evolved solar-like oscillator so far, whose dynamical mass can be derived through a combination of Gaia DR3 astrometry and ground-based SB2 spectroscopy from the Las Cumbres Observatory Global Telescope Network (LCO, Brown et al., 2013). Chaplin et al. (2014) identified and characterized as a solar-like oscillator and the color-magnitude diagram (CMD) suggests a position on the MS or an early subgiant (eSG) (Godoy-Rivera et al., 2025). Molenda-Żakowicz et al. (2013) reported this system as an SB2 system, but did not provide orbital parameters. Eventually, Gaia Collab. et al. (2023) reported this system as ASB1 within the framework of Two-Body Orbit solutions (TBO) of the NSS. In addition to the orbital parameters (=103.750.13 days, =0.390.04), the TBO also reports the inclination angle, , of the orbital plane to be 662 degrees.
@@ 2 Spectroscopic analysis and dynamical masses
We monitored KIC 9693187 with Network of Robotic Echelle Spectrographs (NRES, Brown et al., 2013), mounted on the 1 m telescopes of the northern LCO nodes at McDonald Observatory (USA) and Wise Observatory (Israel). NRES is an échelle spectrograph with a resolution of 53,000. Between April 2023 and June 2024, we have collected 22 spectra of KIC 9693187.
The NRES pipeline calculates the cross-correlation function’s response profile (CCF) from the observed spectra and a synthetic spectrum as a template to compute the radial velocities (RV). The pipeline automatically selected the most appropriate template from the Phoenix model library, based on the Gaia parameters of the target ( = 5700 K, = 4 dex, [Fe/H] = 0 dex). The CCF profiles of the individual spectra, shown in Fig. 4 clearly reveal the SB2 nature of KIC 9693187.
@@ 2.1 Spectral disentangling
To determine the RV semi-amplitudes of the primary and secondary, and , respectively, we applied spectral disentangling (SPD) in Fourier space. This technique optimizes the orbital elements of a binary, including and (Simon and Sturm, 1994; Hadrava, 1995). We note that we refer to the more massive component as the primary, indicated with the subscript 1.
For our SPD analysis of KIC 9693187, we use the FDBinary code (Ilijic et al., 2004). We focused on a 60 nm wide region between 476.1 and 536.1 nm, including the Mg-triplet, which is well suited to determining for both components. The disentangled spectra are depicted in Fig. 4. Except for the orbital period, which we fix to the astrometric solution from the Gaia TBO, we treated all orbital parameters as free parameters. The resulting values are reported in Table 2. The uncertainties were derived from a Gaussian fit to the parameter distributions from 5000 bootstrap simulations (see Pavlovski et al., 2023).
The best solution finds =26.298 0.058 km/s and =29.394 0.066 km/s as the RVs semi-amplitudes for the primary and secondary component. These values reveal a mass ratio of = =0.8947 0.0028 (11% difference in mass).
@@ 2.2 Stellar fundamental parameters
Fundamental stellar parameters were derived through a spectroscopic analysis using an updated version of the 1D/LTE code BACCHUS (Masseron et al., 2016; Hayes et al., 2022). The full analysis and derived parameters are provided in Appendix A.1 and Table 2. We find an effective temperature of 5738 84K and a subsolar metallicity ([Fe/H]=-0.36 0.15dex). The significant enrichment of elements ([/Fe]=+0.25 0.05dex) indicates that the star is member of the older thick-disk population. The projected surface rotation velocity, , of the primary and secondary (6.71.5, and 8.02.0 km/s, respectively) are more than three times the current solar value of 1.96 km/s (Beck and Giles, 2005) and about six times faster than the expected at the MS turnoff of the Sun. Even more puzzling, no photometric signature of spot-modulation is seen in the light curve or PSD (García et al., 2014; Santos et al., 2021, see Fig. 2 below 1 Hz).
@@ 2.3 Dynamical masses
The dynamical masses for both components are given by
From the determined orbital parameters and values taken from spectroscopy (Table 2), along with the and inclination, , provided by Gaia, we obtained =0.9930.046 M⊙ and =0.8890.041 M⊙.
@@ 3.1 Global seismology
To test the asteroseismic scaling relations for the mass (Eq. 2), we used the effective temperature and the global seismic parameters of the oscillating component, obtained with the A2Z pipeline and reported by Mathur et al. (2022). Furthermore, we re-extracted the light-curve and calculated the PSD (for details see Appendix B.1 and Fig. 6) and redetermined the values using the universal pattern module of the apollinaire peakbagging code (Breton et al., 2021), as described in Appendix B.2. The values from both methods agree within the uncertainty of the respective parameters (see Table 3). However, apollinaire leads to smaller uncertainties and a lower value for . Figure 2 compares the results of the scaling relations, without performing a correction on , and following the correction by Mosser et al. (2013) and Li et al. (2023) (see Appendix B.2) with the primary’s dynamical mass.
@@ 3.2 Individual-frequency seismology
Using apollinaire, we extracted 28 frequencies of the spherical degree, = 0, 1, and 2. The extracted frequencies (see Table 5) are depicted in the échelle diagram (Fig. 2). We searched for the best stellar model, constrained by using individual-frequency (IF) seismology. The structural modeling was done with the Modules for Experiments in Stellar Astrophysics stellar evolution code package (MESA, Jermyn et al., 2023; Paxton et al., 2011, and references therein). The frequencies were calculated from each model using the Aarhus adiabatic pulsation codes (Christensen-Dalsgaard, 2008). The modeling setup and input parameters are described in detail in Appendix B.4. The model frequencies in the échelle diagram (Fig. 2) present a good fit. The model that fits best the observational constraints is found with =0.920.01 M⊙ and =1.380.01 R⊙ (see Table 3).
4 Discussion and conclusions
Figure 2 and compares the primary’s dynamical mass with seismically inferred masses from scaling relations and individual frequency modeling (Table 3). From scaling relations, we obtained a mass range of 0.930.03 to 0.980.18 by using the observed from apollinaire and A2Z in the scaling relations for the primary. While the masses from the A2Z values agree within 2%, they suffer from large uncertainties. The apollinaire values lead to a mass that is 6% below, but still within the uncertainties of the dynamical mass.
The large-frequency separation is known to depart from the asymptotic regime increasingly with a decreasing radial order. To correct the obtained masses, we tested two well-established formalisms from the literature (for details see Appendix B.2). The corrections proposed by Mosser et al. (2013) appear to underestimate the seismic mass for this late MS target by 23%, based on the apollinaire set of parameters. The best agreement (2%) is found from the correction of Li et al. (2022).
@@ 4 Discussion and conclusions
The best-fit IF model yields an age of 11.20 0.55 Gyr for the primary component of KIC 9693187 (Table 3). This result confirms the evolved nature of the star, placing it at the main sequence (MS) turnoff, at the point of core hydrogen exhaustion, while the secondary remains on the MS (Fig. 8). This evolutionary configuration is consistent with the system’s -enhanced chemical composition. However, given the primary’s advanced age, the rapid surface rotation remains puzzling.
The uncertainty on the derived dynamical masses remains relatively large compared to the subpercent precision that is in principle achievable with this method (Torres et al., 2010). A major contributor is the relatively large error of on the orbital inclination, which significantly contributes to the uncertainty. Looking ahead, larger sample sizes with improved constraints on orbital inclination and the availability of epoch RV in the forthcoming Gaia DR4, potentially combined with ground-based RV monitoring, will enable rapid progress in refining dynamical mass measurements for solar-like oscillators in binaries.
As suggested by Beck et al. (2024), the combination of Gaia astrometry of non-eclipsing systems with SB2 solutions from ground-based spectroscopy (or from forthcoming Gaia DRs) offers a promising and abundant new source of benchmark calibrators. These developments are particularly timely in light of the upcoming ESA mission PLAnetary Transits and Oscillations of stars (PLATO, Rauer et al., 2025), which will provide high-precision asteroseismic data for ten thousands of solar-like oscillators. A significant number of ASB1 systems identified in Gaia DR3 as astrometric binaries, potentially hosting a solar-like oscillator on the MS, SG, or red-giant phase, have been selected by the authors as part of the Science Calibration and Validation PLATO Input Catalog (scvPIC). The NRES/LCO monitoring presented in this letter is part of a dedicated follow-up program to build an SB2 sample of such high-value targets. This growing number of well-calibrated benchmark stars enables a broader validation of global asteroseismic methods, offering a rich ensemble of well-calibrated stars to test the intricacies of stellar evolution from physics-informed models (e.g., Grossmann et al., 2025; Thomsen et al., 2025; Schimak et al., 2026). Ultimately, these efforts will support PLATO’s core scientific objective of improving the precision of stellar age determinations.